OpenStudy (anonymous):

describe the behavior of y=x^-10 and y=-x^10 a.x->0 b. x->infinity c.x->-infinity (not a less then sign its an arrow)

OpenStudy (loser66):

for y = x ^ -10 = \(\dfrac{1}{x^{10}}\) x \(\rightarrow 0\) \(y\rightarrow \infty\)

OpenStudy (loser66):

x\(\rightarrow \pm\infty \) , y\(\rightarrow 0\)

OpenStudy (loser66):

you can argue on the same way for y = -x^10 don't forget minus sign

OpenStudy (anonymous):

can u explain how u got the answers im kina clueless when it comes to these problems lol

OpenStudy (loser66):